# Generalised Fermat equation: a survey of solved cases

### v1

@@ 2 Equations with special signatures
In this section we will discuss equation (1.2) with special signatures $(p,q,r)$ presented in Table 1.3. We start with the most studied case $(p,q,r)=(4,4,4)$.
@@ 2.1 Equations of signature $(4,4,4)$
for positive integers $a,b,c$. Sections 6.2.2 and 6.3.4 of [75] study equations (2.1) ordered by $a+b+c$, decide the existence of a solution $(x,y,z)\neq(0,0,0)$ for all equations with $a+b+c\leq 62$, and solves Problem 1.7 for the equations with $a+b+c<39$. The first equation (2.1) for which Problem 1.7 is left open in [75] is
@@
However, there are equations (2.1) with $a+b+c<39$ that have easy-to-find non-trivial solutions, which solves Problem 1.7 but leaves open Problem 1.6. For example, this happens if $a+b=c$, in which case $(x,y,z)=(\pm 1,\pm 1,\pm 1)$ are the solutions. The smallest non-trivial example in this category is the equation
@@
whose only primitive solutions are $(x,y,z)=(\pm 1,\pm 1,\pm 1)$, as proved by Jeremy Rouse on Mathoverflow website [^6] https://mathoverflow.net/questions/480114. Other non-trivial examples with $a+b=c$ are the equations
$(a)\,\,x^{4}+3y^{4}=4z^{4},\quad\text{and}\quad(b)\,\,x^{4}+8y^{4}=9z^{4},$ (2.2)
Lucas [57, pp. 630] proved that these equations have no primitive integer solutions other than (±$(\pm 1,\pm 1,\pm 1)$.
In case $a=b=1$, equation (2.1) reduces to
@@
Problem 1.7 for this equation reduces to the question whether a positive integer $c$ can be represented as a sum of two rational fourth powers. Cohen [44, Section 6.6] solved this problem for all integers $c$ in the range $0001\leq c\leq 10,$ 000, except of $c=4481$, $7537$ and $8882$. These values of $c$ have been addressed in [24] and [25], hence Problem 1.7 for (2.3) is now solved for all $0001\leq c\leq 10,$ 000. In particular, $c=5906=(25/17)^{4}+(149/17)^{4}$ is the smallest integer that is the sum of two fourth powers of rational numbers but not the sum of two fourth powers of integers.
Grigorov and Rizov [77] proved that if $c>2$ is a fourth-power free integer and the rank of elliptic curve $v^{2}=u^{3}-cu$ is $1$, then the equation (2.3) has no non-zero integer solutions.
For values of $c$ when (2.3) is solvable, such as $c=17$, the question to find all primitive solutions (that is, Problem 1.6) has not been addressed in [44, Section 6.6]. The mentioned case $c=17$ has been solved in 2001 by Flynn and Wetherell [67].
@@
The primitive positive integer solutions to the equation (2.3) with $c=17$, that is,
@@
A combination of Theorem 2.1 with elementary methods discussed in the introduction solves equation (2.3) for all $1\leq c\leq 81$, and in fact for all $1\leq c\leq 100$ except of $c=82$ and $c=97$. In 2023, using the methods developed to prove Theorem 2.1 with some extra tweaks, Pădurariu [113] was able to resolve the case $c=97$.
@@
The primitive positive integer solutions to the equation (2.3) with $c=97$, that is,
@@
Equation (2.3) with $z\neq 0$ can be rewritten as $(x/z)^{4}+(y/z)^{4}=c$, and reduces to finding a representation of $c$ as a sum of two rational fourth powers. A similar problem of representing an integer $b$ as a difference of two fourth powers reduces to equation
@@
with $y\neq 0$. The problem of existence of integer solutions to (2.4) has been solved in [75, Section 6.3.4] for $1\leq b\leq 218$, and is left open for $b=219$. However, if a non-trivial solution to (2.4) exists, the problem of finding all solutions has not been studied in [75]. In 2020, Söderlund [133] resolved this problem for $b=34$.
@@
Taclay and Bacani [137] consider equation (2.4) with $b=2p$ where $p$ is a prime and they prove the following result.
@@
The equation (2.1) with coefficients $a,b,c$ being perfect cubes was studied in [102, 103], and we have the following result.
@@
Examples of values $(s,t,u,n,v,w)$ satisfying these conditions are $(7,15,23,8280,4991,13335)$,
$(7,15,23,8280,16583,15855)$ and $(7,15,23,11040,3703,14175)$ which correspond to equations
$12176^{3}x^{4}+6473^{3}y^{4}=3881^{3}z^{4},\quad 20512^{3}x^{4}+353^{3}y^{4}=5297^{3}z^{4},\quad 18208^{3}x^{4}+10313^{3}y^{4}=6073^{3}z^{4},$
@@ 2.2 Equations of signature $(2,4,r)$
Let us first consider the equation (2.6) with $a=2$ and $b=c=1$, that is,
@@
In 2008, Dieulefait and Urroz [61] proved that if $r>349$ is a prime, then equation (2.7) does not have non-trivial primitive solutions. In 2010, Bennett, Ellenberg and Ng [10] significantly strengthened this result and completely solved equation (2.7) for all integers $r\geq 5$.
@@
The equation (2.7) has no non-trivial primitive solutions for integer $r\geq 6$, while for $r=5$ its only non-trivial primitive integer solutions are $(x,y,z)=(\pm 11,\pm 1,3)$.
Let us now consider the equation (2.6) with $a=3$ and $b=c=1$. Dieulefait and Urroz [61] proved that if $r>131$ is a prime, then the equation
@@
In 2022, Pacetti and Villagra Torcomian [111] studied the equation (2.6) with $a\in\{5,6,7\}$, $b=c=1$ and prime $p$, that is,
@@
Let $(a,p_{0})\in\{(5,499),(6,563),(7,349)\}$. Assume that $p>p_{0}$ is prime. Then the equation (2.8) has no non-trivial primitive solutions.
@@
Let $s>3$ be a prime and let $s\equiv 3$ mod 8 and $s\neq 2t^{2}+1$, or $s\equiv 5$ mod 8 and $s\neq t^{2}+4$. In addition, let $p$ be a prime satisfying $p>(8\sqrt{s+1}+1)^{16(s-1)}$. Then equation (2.9) has no non-trivial primitive integer solutions.
Examples of $s$ satisfying the conditions of Theorem 2.8 are $s=11,37,43,\dots$.
@@
Equation (2.11) has been further studied by Bennett [4], who proved a version of Theorem 2.10 with a smaller set of excluded primes.
@@
Let $s$ be a prime not of the form $s=2^{2^{k}}+1$ for integer $k\geq 1$, and let $\alpha,\beta$ be non-negative integers. Let $p$ be a prime satisfying $p>(\sqrt{8s+8}+1)^{2s-2}$. Then equation (2.11) has no non-trivial primitive solutions.
Primes of the form $s=2^{2^{k}}+1$ are called Fermat primes. It is widely believed that the only such primes are $s=3,5,17,257$ and $65537$. The case $s=3$ corresponds to $k=0$, while Theorem 2.11 fails to apply only if $k\geq 1$. Bennett [4] also solved equation (2.11) for the exceptional cases $s=5$ and $17$, at least for $(\alpha,\beta)=(0,1)$.
@@
If $s$ is a prime with $502\leq s<$ 50, $s\neq 5$, $s\neq 17$, $\alpha$ and $\beta$ are non-negative integers, and $p>5$ is a prime, then equation (2.11) has no non-trivial primitive integer solutions.
@@
Let $p>19$ be a prime number such that $p\neq 97$ and $p\equiv 1,3$ modulo $8$. Then, equation (2.12) with $b=6$ has the only non-trivial primitive solutions $(x,y,z)=(1,\pm 20,\pm 7)$.
@@
then equation (2.12) has no non-trivial primitive solutions.
@@
Let $p>900$ be a prime number. Then, there are no non-trivial primitive solutions of the equation (2.12) with $b=129$.
Cao [34] studied the equation (2.12) with $b=p$, that is,
@@
and proved that if $p\equiv 1$ modulo $4$ is a prime and [^7] Here, $B_{n}$ denotes the $n$-th Bernoulli number. $p\nmid B_{(p-1)/2}$, then the equation (2.13) has no integer solutions with $\gcd(y,z)=1$, $p|y$ and $2|z$.
@@
where $h(-a)$ denotes the class number [^8] A list of the class numbers for the first 10,000 square-free $a$ can be found at https://oeis.org/A000924/b000924.txt. of $\mathbb{Q}(\sqrt{-a})$, equation (2.14) has no non-trivial primitive integer solutions.
@@ 2.3 Equations of signature $(2,6,r)$
Chen [41] studied equation (1.2) with $p$ prime, $q=6$, $b=r=2$ and $a=c=1$, that is,
@@
Let $p$ be a prime such that $p\equiv 1,7$ (mod $24$) and $p\neq 7$. Then the equation (2.15) does not have any non-trivial primitive integer solutions except those with $x=\pm 1$.
The equation (1.2) with $r$ prime, $q=6$, $p=2$ and $a=c=1$, that is,
@@
Let $r>257$ be a prime number. Then equation (2.17) has no non-trivial primitive solutions.
@@
Let $r\geq 3$ be an integer. If $r\neq 4$, then equation (2.18) has no non-trivial primitive integer solutions, while for $r=4$ its only non-trivial primitive integer solutions are $(x,y,z)=(\pm 47,\pm 2,\pm 7)$.
The equation (2.16) with $b=6$, that is,
@@
Let $r>563$ be a prime number. Then, equation (2.19) has no non-trivial primitive solutions.
Equation (2.16) with some other values of $b$ has been studied in [95], where the following theorem has been proven.
@@
Let $(b,r_{0})$ be as in Table 2.1. Assume that $r\geq r_{0}$ is a prime satisfying the congruence conditions in Table 2.1. Then equation (2.16) has no non-trivial primitive integer solutions.
@@
$5$ $1033$
$7$ $337$ $r\equiv 5,7$ (mod $12$)
$11$ $557$ $r\equiv 3$ (mod 4)
$13$ $3491$
$15$ $743$ $r\equiv 5,7,15,17,$ or $19$ (mod 24)
$19$ $1031$
Table 2.1: Details for Theorem 2.22.
@@ 2.4 Equations of signature $(4,4,3)$
Let us now discuss the equation (1.2) with $(p,q,r)=(4,4,3)$ and $a=c=1$, that is, equation
@@ 2.5 A systematic approach
Monograph [75] suggests studying Diophantine equations systematically. It defines the size of equation (1.2) as $H=|a|2^{p}+|b|2^{q}+|c|2^{r}$, and lists all non-trivial equations of size $H\leq 60$. Equations of small size with $|abc|=1$ are covered in Table 1.1. Equations $x^{4}+y^{4}=2z^{3}$ and $y^{3}+z^{3}=2x^{4}$ of size $H=48$ are special cases of Theorems 4.32 and 4.3, respectively. Equation $2x^{2}+y^{4}=z^{5}$ of size $H=56$ is the $r=5$ case of Theorem 2.6. Equation $x^{4}+2y^{4}=z^{3}$ of the same size has been solved in Theorem 2.23. Equations $x^{4}+2y^{3}+2z^{3}=0$, $x^{4}-y^{4}=2z^{3}$ and $x^{4}-y^{4}=3z^{3}$ of sizes $H=48$, $48$ and $56$, respectively, have been solved in [143, Appendix A]. All other non-trivial equations of size $H\leq 60$ are listed in Table 2.2. All these equations except $x^{5}+2y^{3}+z^{3}=0$ have special signatures, and we were not able to find a reference where any of them has been solved. We invite the reader to investigate these equations.
@@
$40$ $x^{4}+2y^{3}+z^{3}=0$ $56$ $02x^{4}-y^{4}+z^{3}=$ 0 $60$ $x^{5}+y^{4}+3z^{2}=0$
$48$ $x^{4}+3y^{3}+z^{3}=0$ $56$ $02x^{4}+2y^{3}+z^{3}=$ 0 $60$ $x^{5}+y^{4}-3z^{2}=0$
$56$ $x^{4}+3y^{3}+2z^{3}=0$ $56$ $x^{5}+2y^{3}+z^{3}=0$
$56$ $x^{4}+4y^{3}+z^{3}=0$ $56$ $x^{5}+y^{4}-2z^{2}=0$
Table 2.2: Uninvestigated equations (1.2) with $|a|2^{p}+|b|2^{q}+|c|2^{r}\leq 60$.

### v2

@@ 2 Equations with special signatures
In this section we will discuss Eq. (1.1) with special signatures $(p,q,r)$ presented in Table 1.3. We start with the most studied case $(p,q,r)=(4,4,4)$.
@@ 2.1 Equations of signature $(4,4,4)$
for positive integers $a,b,c$. Sections 6.2.2 and 6.3.4 of [77] and Sections 6.4 and 6.5 of [148] study Eqs. (2.1) ordered by $a+b+c$, decide the existence of a solution $(x,y,z)\neq(0,0,0)$ for all equations with $a+b+c\leq 62$, and solves Problem 1.8 for the equations with $a+b+c<39$. The first Eq. (2.1) for which Problem 1.8 is left open in [77] is
@@
However, there are Eqs. (2.1) with $a+b+c<39$ that have easy-to-find non-trivial solutions, which solves Problem 1.8 but leaves open Problem 1.7. For example, this happens if $a+b=c$, in which case $(x,y,z)=(\pm 1,\pm 1,\pm 1)$ are the solutions. The smallest non-trivial example in this category is the equation
@@
whose only primitive solutions are $(x,y,z)=(\pm 1,\pm 1,\pm 1)$, as proved by Jeremy Rouse on Mathoverflow website. [^9] https://mathoverflow.net/questions/480114 Other non-trivial examples with $a+b=c$ are the equations
$(a)\,\,x^{4}+3y^{4}=4z^{4},\quad\text{and}\quad(b)\,\,x^{4}+8y^{4}=9z^{4}.$
Lucas (see [59, pp. 630]) proved that these equations have no primitive integer solutions other than (±$(\pm 1,\pm 1,\pm 1)$.
In the case $a=b=1$, Eq. (2.1) reduces to
@@
Problem 1.8 for this equation reduces to the question whether a positive integer $c$ can be represented as a sum of two rational fourth powers. Cohen [46, Section 6.6] solved this problem for all integers $c$ in the range $0001\leq c\leq 10,$ 000, except of $c=4481$, $7537$ and $8882$. These values of $c$ have been addressed in [25, 26], hence Problem 1.8 for (2.2) is now solved for all $0001\leq c\leq 10,$ 000. In particular, $c=5906=(25/17)^{4}+(149/17)^{4}$ is the smallest integer that is the sum of two fourth powers of rational numbers but not the sum of two fourth powers of integers.
Grigorov and Rizov [79] proved that if $c>2$ is a fourth-power free integer and the rank of elliptic curve $v^{2}=u^{3}-cu$ is $1$, then Eq. (2.2) has no non-zero integer solutions.
For values of $c$ when (2.2) is solvable, such as $c=17$, the question to find all primitive solutions (that is, Problem 1.7) has not been addressed in [46, Section 6.6]. The mentioned case $c=17$ has been solved in 2001 by Flynn and Wetherell [69].
@@
The primitive positive integer solutions to Eq. (2.2) with $c=17$, that is,
@@
A combination of Theorem 2.1 with elementary methods discussed in the introduction solves Eq. (2.2) for all $1\leq c\leq 81$, and in fact for all $1\leq c\leq 100$ except of $c=82$ and $c=97$. In 2023, using the methods developed to prove Theorem 2.1 with some extra tweaks, Pădurariu [116] was able to resolve the case $c=97$.
@@
The primitive positive integer solutions to Eq. (2.2) with $c=97$, that is,
@@
Eq. (2.2) with $z\neq 0$ can be rewritten as $(x/z)^{4}+(y/z)^{4}=c$, and reduces to finding a representation of $c$ as a sum of two rational fourth powers. A similar problem of representing an integer $b$ as a difference of two fourth powers reduces to equation
@@
with $y\neq 0$. The problem of existence of integer solutions to (2.3) has been solved in [77, Section 6.3.4] for $1\leq b\leq 218$, and is left open for $b=219$. However, if a non-trivial solution to (2.3) exists, the problem of finding all solutions has not been studied in [77]. In 2020, Söderlund [138] resolved this problem for $b=34$.
@@
Taclay and Bacani [142] consider Eq. (2.3) with $b=2p$ where $p$ is a prime and they prove the following result.
@@
The smallest prime $p$ for which Eq. (2.4) is not covered by Theorems 2.3 and 2.4 is $p=97$.
Eq. (2.1) with coefficients $a,b,c$ being perfect cubes was studied in [105, 106], and we have the following result.
@@
Examples of values $(s,t,u,n,v,w)$ satisfying these conditions are $(7,15,23,\\ 8280,4991,13335)$, $(7,15,23,8280,16583,15855)$ and $(7,15,23,11040,3703,\\ 14175)$ which correspond to equations
$12176^{3}x^{4}+6473^{3}y^{4}=3881^{3}z^{4},\quad 20512^{3}x^{4}+353^{3}y^{4}=5297^{3}z^{4},$
$18208^{3}x^{4}+10313^{3}y^{4}=6073^{3}z^{4},$
@@ 2.2 Equations of signature $(2,4,r)$
Let us first consider Eq. (2.5) with $a=2$ and $b=c=1$, that is,
@@
In 2008, Dieulefait and Urroz [65] proved that if $r>349$ is a prime, then Eq. (2.6) does not have non-trivial primitive solutions. In 2010, Bennett, Ellenberg and Ng [11] significantly strengthened this result and completely solved Eq. (2.6) for all integers $r\geq 5$.
@@
Eq. (2.6) has no non-trivial primitive solutions for integer $r\geq 6$, while for $r=5$ its only non-trivial primitive integer solutions are $(x,y,z)=(\pm 11,\pm 1,3)$.
Let us now consider Eq. (2.5) with $a=3$ and $b=c=1$. Dieulefait and Urroz [65] proved that if $r>131$ is a prime, then the equation
@@
In 2022, Pacetti and Villagra Torcomian [114] studied Eq. (2.5) with $a\in\{5,6,7\}$, $b=c=1$ and prime $p$, that is,
@@
Let $(a,p_{0})\in\{(5,499),(6,563),(7,349)\}$. Assume that $p>p_{0}$ is prime. Then Eq. (2.7) has no non-trivial primitive solutions.
@@
Let $s>3$ be a prime and let $s\equiv 3$ mod 8 and $s\neq 2t^{2}+1$, or $s\equiv 5$ mod 8 and $s\neq t^{2}+4$. In addition, let $p$ be a prime satisfying $p>(8\sqrt{s+1}+1)^{16(s-1)}$. Then Eq. (2.8) has no non-trivial primitive integer solutions.
Examples of $s$ satisfying the conditions of Theorem 2.8 are $s=11$, $37$, $43$ and so on.
@@
Eq. (2.10) has been further studied by Bennett [5], who proved a version of Theorem 2.10 with a smaller set of excluded primes.
@@
Let $s$ be a prime not of the form $s=2^{2^{k}}+1$ for integer $k\geq 1$, and let $\alpha,\beta$ be non-negative integers. Let $p$ be a prime satisfying $p>(\sqrt{8s+8}+1)^{2s-2}$. Then Eq. (2.10) has no non-trivial primitive integer solutions.
Primes of the form $s=2^{2^{k}}+1$ are called Fermat primes. It is widely believed that the only such primes are $s=3,5,17,257$ and $65537$. The case $s=3$ corresponds to $k=0$, while Theorem 2.11 fails to apply only if $k\geq 1$. Bennett [5] also solved Eq. (2.10) for the exceptional cases $s=5$ and $17$, at least for $(\alpha,\beta)=(0,1)$.
@@
If $s$ is a prime with $502\leq s<$ 50, $s\neq 5$, $s\neq 17$, $\alpha$ and $\beta$ are non-negative integers, and $p>5$ is a prime, then Eq. (2.10) has no non-trivial primitive integer solutions.
@@
Let $p>19$ be a prime number such that $p\neq 97$ and $p\equiv 1,3$ modulo $8$. Then, Eq. (2.11) with $b=6$ has the only non-trivial primitive integer solutions $(x,y,z)=(1,\pm 20,\pm 7)$.
@@
then Eq. (2.11) has no non-trivial primitive solutions.
@@
Let $p>900$ be a prime number. Then, there are no non-trivial primitive solutions of Eq. (2.11) with $b=129$.
Cao [35] studied Eq. (2.11) with $b=p$, that is,
@@
and proved that if $p\equiv 1$ modulo $4$ is a prime and [^10] Here, $B_{n}$ denotes the $n$th Bernoulli number. $p\nmid B_{(p-1)/2}$, then Eq. (2.12) has no integer solutions with $\gcd(y,z)=1$, $p|y$ and $2|z$.
@@
where $h(-a)$ denotes the class number [^11] A list of the class numbers for the first 10,000 square-free $a$ can be found at https://oeis.org/A000924/b000924.txt. of $\mathbb{Q}(\sqrt{-a})$, Eq. (2.13) has no non-trivial primitive integer solutions.
@@ 2.3 Equations of signature $(2,6,r)$
Chen [42] studied Eq. (1.1) with $p$ prime, $q=6$, $b=r=2$ and $a=c=1$, that is,
@@
Let $p$ be a prime such that $p\equiv 1,7$ (mod $24$) and $p\neq 7$. Then Eq. (2.14) does not have any non-trivial primitive integer solutions except those with $x=\pm 1$.
Eq. (1.1) with $r$ prime, $q=6$, $p=2$ and $a=c=1$, that is,
@@
Let $r>257$ be a prime number. Then Eq. (2.16) has no non-trivial primitive integer solutions.
@@
Let $r\geq 3$ be an integer. If $r\neq 4$, then Eq. (2.17) has no non-trivial primitive integer solutions, while for $r=4$ its only non-trivial primitive integer solutions are $(x,y,z)=(\pm 47,\pm 2,\pm 7)$.
Eq. (2.15) with $b=6$, that is,
@@
Let $r>563$ be a prime number. Then, Eq. (2.18) has no non-trivial primitive integer solutions.
Eq. (2.15) with some other values of $b$ has been studied in [97], where the following theorem has been proven.
@@
Let $(b,r_{0})$ be as in Table 2.1. Assume that $r\geq r_{0}$ is a prime satisfying the congruence conditions in Table 2.1. Then Eq. (2.15) has no non-trivial primitive integer solutions.
Table 2.1: Details for Theorem 2.22.
@@
5 1033
7 337 $r\equiv 5,7$ (mod 12)
11 557 $r\equiv 3$ (mod 4)
13 3491
15 743 $r\equiv 5,7,15,17,$ or 19 (mod 24)
19 1031
@@ 2.4 Equations of signature $(4,4,3)$
Let us now discuss Eq. (1.1) with $(p,q,r)=(4,4,3)$ and $a=c=1$, that is, equation
@@ 2.5 A systematic approach
Table 2.2: Uninvestigated equations (1.1) with $H=|a|2^{p}+|b|2^{q}+|c|2^{r}\leq 60$.
@@
40 $x^{4}+2y^{3}+z^{3}=0$ 56 $x^{4}+4y^{3}+z^{3}=0$ 56 $x^{5}+y^{4}-2z^{2}=0$
48 $x^{4}+3y^{3}+z^{3}=0$ 56 $02x^{4}-y^{4}+z^{3}=$ 0 60 $x^{5}+y^{4}+3z^{2}=0$
56 $x^{4}+3y^{3}+2z^{3}=0$ 56 $x^{5}+2y^{3}+z^{3}=0$ 60 $x^{5}+y^{4}-3z^{2}=0$
Monograph [77] suggests studying Diophantine equations systematically. It defines the size of Eq. (1.1) as $H=|a|2^{p}+|b|2^{q}+|c|2^{r}$, and lists all non-trivial equations of size $H\leq 60$. Equations of small size with $|abc|=1$ are covered in Table 1.1. Equations $x^{4}+y^{4}=2z^{3}$ and $y^{3}+z^{3}=2x^{4}$ of size $H=48$ are special cases of Theorems 4.32 and 4.3, respectively. Equation $2x^{2}+y^{4}=z^{5}$ of size $H=56$ is the $r=5$ case of Theorem 2.6. Equation $x^{4}+2y^{4}=z^{3}$ of the same size has been solved in Theorem 2.23. Equations $x^{4}+2y^{3}+2z^{3}=0$, $x^{4}-y^{4}=2z^{3}$ and $x^{4}-y^{4}=3z^{3}$ of sizes $H=48$, $48$ and $56$, respectively, have been solved in [148, Appendix A]. All other non-trivial equations of size $H\leq 60$ are listed in Table 2.2. All these equations except $x^{5}+2y^{3}+z^{3}=0$ have special signatures, and we were not able to find a reference where any of them has been solved. We invite the reader to investigate these equations.